At what distance does a man 11/2ft in height, subtend an angle of 15"?
 

Grade:11

1 Answers

Avinash
5 years ago
Assume triangle with height 11/2 i.e 5.5 which subtends an angle 15°at a point at a distance of X from that point So tan15°=5.5/X2-√3=5.5/Xtherefore X=5.5/2-√3 X=5.5(2+√3)/4-3 X=5.5(2+√3)